# references.bib

### v1

@@ Abstract
Autonomous Vehicles (AVs) collect and pseudo-label terabytes of multi-modal data localized to HD maps during normal fleet testing. However, identifying interesting and safety-critical scenarios from uncurated driving logs remains a significant challenge. Traditional scenario mining techniques are error-prone and prohibitively time-consuming, often relying on hand-crafted structured queries. In this work, we revisit spatio-temporal scenario mining through the lens of recent vision-language models (VLMs) to detect whether a described scenario occurs in a driving log and, if so, precisely localize it in both time and space. To address this problem, we introduce RefAV, a large-scale dataset of $10,000$ diverse natural language queries that describe complex multi-agent interactions relevant to motion planning derived from $1000$ driving logs in the Argoverse 2 Sensor dataset. We evaluate several referential multi-object trackers and present an empirical analysis of our baselines. Notably, we find that naively repurposing off-the-shelf VLMs yields poor performance, suggesting that scenario mining presents unique challenges. Our code and dataset are available on Github and Argoverse.
@@ 1 Introduction
Contributions. We present three major contributions. First, we introduce RefAV, a large-scale dataset designed to evaluate VLMs on 3D scene understanding and spatio-temporal localization. Our extensive experiments highlight the limitations of current methods, and demonstrate the effectiveness of our proposed program synthesis-based approach. Lastly, we host a challenge at CVPR 2025 in conjunction with the Workshop on Autonomous Driving to encourage broad community involvement in addressing this challenging problem.
@@ 3.1 Creating RefAV
LLM-based Procedural Scenario Generation. Our key insight is that many complex scenarios (e.g. find all cars accelerating while changing lanes) can be broken down into simpler atomic actions (e.g. find cars, accelerating, and changing lanes), and a large set of such atomic actions can be composed to generate new diverse scenarios. To this end, we first define 28 atomic functions based on nuPlan’s list of planning scenarios (cf. Figure 2 left). We include a full API listing in Appendix F. To generate a new scenario definition, we provide an LLM (e.g. Claude 3.5 Sonnet) with the full API listing, along with in-context examples of real compositions. We prompt the LLM to generate permutations of the atomic functions and describe the generated code with a natural language description (cf. Figure 2 middle). We execute the generated code to filter ground-truth tracks from all $1000$ logs to identify true positive matches. We aggregate all true positive log-prompt matches (>50K) and sample 8,000 true positive log-prompt pairs that maximize scenario diversity. Lastly, we randomly sample 2,000 true negative log-prompt pairs.
@@ 4.4 CVPR 2025 Scenario Mining Challenge
We are hosting a challenge at CVPR 2025 to encourage broad community involvement in addressing spatio-temporal scenario mining. To incentivize participation, top teams can win cash prizes. Although our competition will continue until June 8th 2025, we have already received submissions from three teams (some submissions are private).
@@ 4.5 Limitations and Future Work
Figure 4: Examples of Multi-Agent Interactions. We visualize representative examples from RefAV to highlight the diversity of our dataset. In (a), we capture the interactions between vulnerable road users and vehicles at a crowded intersection. Scenario (b) presents an atypical instance of a common multi-agent interaction (e.g. pedestrian walking a dog). In (c), we show a complex ego-vehicle trajectory that involves multiple moving vehicles. Scenario (d) illustrates an example of a rare multi-object interaction. In (e), we highlight a scenario that might requires evasive maneuvers from the ego-vehicle (e.g. the occluded pedestrian might cross the path of the ego-vehicle). Finally, subfigure (f) visualizes an example scene graph derived from AV2 annotations, where edges represent links between the referred object and related objects in the environment. Note that we show referred objects in green, related objects in blue, and all other objects in red.
@@ 6 Acknowledgments
This work was supported in part by funding the NSF GRFP (Grant No. DGE2140739).
@@ Appendix A Implementation Details
We include the API listing in Appendix F. We found this prompt consistently produces reasonable code across all tested LLMs. After running the synthesized programs, we post-process the output tracks in accordance with dataset practices. We remove all length one tracks as noise and discard object relationships where the spatial distance exceeds 50 meters.
@@ Appendix E More Dataset Statistics
Appendix F API Listing

### v2

@@ Abstract
Autonomous Vehicles (AVs) collect and pseudo-label terabytes of multi-modal data localized to HD maps during normal fleet testing. However, identifying interesting and safety-critical scenarios from uncurated driving logs remains a significant challenge. Traditional scenario mining techniques are error-prone and prohibitively time-consuming, often relying on hand-crafted structured queries. In this work, we revisit spatio-temporal scenario mining through the lens of recent vision-language models (VLMs) to detect whether a described scenario occurs in a driving log and, if so, precisely localize it in both time and space. To address this problem, we introduce RefAV, a large-scale dataset of $10,000$ diverse natural language queries that describe complex multi-agent interactions relevant to motion planning derived from $1000$ driving logs in the Argoverse 2 Sensor dataset. We evaluate several referential multi-object trackers and present an empirical analysis of our baselines. Notably, we find that naively repurposing off-the-shelf VLMs yields poor performance, suggesting that scenario mining presents unique challenges. Lastly, we discuss our recent CVPR 2025 competition and share insights from the community. Our code and dataset are available on Github and Argoverse.
@@ 1 Introduction
Contributions. We present three major contributions. First, we introduce RefAV, a large-scale dataset designed to evaluate VLMs on 3D scene understanding and spatio-temporal localization. Our extensive experiments highlight the limitations of current methods, and demonstrate the effectiveness of our proposed program synthesis-based approach. Lastly, we highlight the results of our recent CVPR 2025 challenge hosted in conjunction with the Workshop on Autonomous Driving.
@@ 3.1 Creating RefAV
LLM-based Procedural Scenario Generation. Our key insight is that many complex scenarios (e.g. find all cars accelerating while changing lanes) can be broken down into simpler atomic actions (e.g. find cars, accelerating, and changing lanes), and a large set of such atomic actions can be composed to generate new diverse scenarios. To this end, we first define 28 atomic functions based on nuPlan’s list of planning scenarios (cf. Figure 2 left). We include a full API listing in Appendix G. To generate a new scenario definition, we provide an LLM (e.g. Claude 3.5 Sonnet) with the full API listing, along with in-context examples of real compositions. We prompt the LLM to generate permutations of the atomic functions and describe the generated code with a natural language description (cf. Figure 2 middle). We execute the generated code to filter ground-truth tracks from all $1000$ logs to identify true positive matches. We aggregate all true positive log-prompt matches (>50K) and sample 8,000 true positive log-prompt pairs that maximize scenario diversity. Lastly, we randomly sample 2,000 true negative log-prompt pairs.
@@ 4.4 CVPR 2025 Scenario Mining Challenge
We hosted a challenge at CVPR 2025 to encourage broad community involvement in addressing spatio-temporal scenario mining. Our competition recieved submissions from eight teams (some submissions are private) at the close of our competition on June 8th, 2025 AOE. Notably, four teams beat our best baseline. We present the current top three teams at the bottom of Table 2. We summarize the contributions of the top three teams in Appendix F and include a link to full technical reports and code here.
@@ 4.5 Limitations and Future Work
Figure 4: Examples of Multi-Agent Interactions. We visualize representative examples from RefAV to highlight the diversity of our dataset. In (a), we capture the interactions between vulnerable road users and vehicles at a crowded intersection. Scenario (b) presents an atypical instance of a common multi-agent interaction (e.g. pedestrian walking a dog). In (c), we show a complex ego-vehicle trajectory that involves multiple moving vehicles. Scenario (d) illustrates an example of a rare multi-object interaction. In (e), we highlight a scenario that might require evasive maneuvers from the ego-vehicle (e.g. the occluded pedestrian might cross the path of the ego-vehicle). Finally, subfigure (f) visualizes a scenario with a multiple-step relationship (e.g. the official signaler is standing inside of a construction zone). Note that we show referred objects in green, related objects in blue, and all other objects in red.
@@ 6 Acknowledgments
This work was supported in part by funding from Bosch Research and the NSF GRFP (Grant No. DGE2140739). We would also like to thank Ishan Khatri for his feedback on early drafts
@@ Appendix A Implementation Details
We include the API listing in Appendix G. We found this prompt consistently produces reasonable code across all tested LLMs. After running the synthesized programs, we post-process the output tracks in accordance with dataset practices. We remove all length one tracks as noise and discard object relationships where the spatial distance exceeds 50 meters.
@@ Appendix E More Dataset Statistics
Appendix F Summary of CVPR 2025 Competition Top Performers
We summarize the contributions of top teams below. We present full technical reports and code here.
Zeekr UMCV extends RefProg with global context-aware generation to processes all queries collectively for improved reasoning and consistency. They also propose a multi-agent refinement system where a secondary refiner agent iteratively debugs and enhances the generated code.
Mi3 UCM proposes two key improvements to RefProg: (1) Fault-Tolerant Iterative Code Generation (FT-ICG), which refines faulty code by re-prompting the LLM with error feedback until successful execution, and (2) Enhanced Prompting for Spatial Relational Functions (EP-SRF), which clarifies function semantics to prevent misinterpretations of spatial relationships between objects.
ZXH refines RefProg by prompting LLMs with a batch of queries and in-context examples instead of prompting for each query one at a time. Importantly, ZXH’s approach relies on in-context learning and batch prompting to improve consistency and efficiency.
Appendix G API Listing